OpenStudy (anonymous):

Evaluate integral from 0 to Pi/2 sin(0) d0

OpenStudy (anonymous):

\[\int\limits_{0}^{\Pi/2}\sin(0) d0\]

OpenStudy (anonymous):

[-cosx]0to pi/2 [0=1]=1

OpenStudy (anonymous):

thanks a lot!

OpenStudy (anonymous):

\[\int\limits_{-2}^{3}\sin ^{2}(x) + \cos ^{2}(x) dx\]